# {textblock}

### v1

@@ 1 Introduction
In this paper, the existing FBS from [9] approach is extended to incorporate specific constraints, resulting in the conditional filter band selection (CFBS) algorithm. While the multispectral classification performance is maximized, the number of filters used is kept as small as possible. Furthermore, the influence of noise can be estimated and limited by making use of this novel approach.
@@ 2 Related Work
To address this problem, Sippel et al. developed the fast binary search [9], which is designed to significantly reduce the computational cost compared to the fill search, while still providing an optimal filter selection. It is a common method to find a near optimum solution for selecting $N$ out of $K$ filters from an available filter set, while the spectral behavior of the cameras and lenses are assumed to be perfect.
@@ 3 Proposed Conditional Filter Band Selection Algorithm
So far when performing a filter selection, typically a fixed number of filters to select is given as a constraint. However, sometimes it is advantageous to chose a minimum number of filters or to add further restrictions and conditions. Therefore, the result of the FBS algorithm will be further processed in the conditional filter band selection approach. Our CFBS method performs a systematic search over the optimal filter selection result ŝ and returns the minimum amount $\textbf{{s}}_{\min}$ of filters needed, while still guaranteeing a good multispectral classification of the material
ŝ$_{\min}=\text{CFBS}(\textbf{{s}})~{}~{}~{}.$ (6)

### v2

@@ 1 Introduction
In this paper, the existing FBS approach from [9] is extended to incorporate specific constraints, resulting in the conditional filter band selection (CFBS) algorithm. While the multispectral classification performance is maximized, the number of filters used is kept as small as possible. Furthermore, the influence of noise can be estimated and limited by making use of this novel approach.
@@ 2 Related Work
To address this problem, Sippel et al. developed the fast binary search [9], which is designed to significantly reduce the computational cost compared to the full search, while still providing an optimal filter selection. It is a common method to find a near optimum solution for selecting $N$ out of $K$ filters from an available filter set, while the spectral behavior of the cameras and lenses are assumed to be perfect.
@@ 3 Proposed Conditional Filter Band Selection Algorithm
So far when performing a filter selection, typically a fixed number of filters to select is given as a constraint. However, sometimes it is advantageous to chose a minimum number of filters or to add further restrictions and conditions. Therefore, the result of the FBS algorithm will be further processed in the conditional filter band selection approach. Our CFBS method performs a systematic search over the optimal filter selection result s and returns the minimum amount $\textbf{{s}}_{\min}$ of filters needed, while still guaranteeing a good multispectral classification of the material
$\vspace{-0.2cm}\textbf{{s}}_{\min}=\text{CFBS}(\textbf{{s}})~{}~{}~{}.$ (6)